# Co-Seg++: Mutual Prompt-Guided Collaborative Learning for Versatile Medical Segmentation

### v1

@@ 4.1.1 Datasets
PUMA [16] is a melanoma nuclei and tissue segmentation dataset containing 206 pathological slides scanned at 40$\times$ magnification with a resolution of $1024\times 1024$. The dataset covers two tissue regions (i.e., tumor and stroma) and three nuclei types (i.e., tumor, lymphocyte, and other cells).
GlaS [17] dataset, derived from the 2015 Gland Segmentation Challenge, comprises 165 histopathology images captured at 20$\times$ magnification with glandular regions and epithelial, connective and lympho-reticular nuclei instances. The image size varies from $574\times 433$ to $775\times 522$.
@@ 4.1.2 Implementation Details
We perform all experiments on a NVIDIA A5000 GPU using PyTorch. For fair comparisons, all semantic and instance segmentation methods are implemented with the same training settings and configurations. We utilize the pre-trained SAM ViT-H [32] structure as the image encoder of medical SAM baselines. We adopt the Hirea-L structure [43] as the image encoder of our Co-Seg++. We apply the optimizer using Adam with an initial learning rate of $1\times 10^{-4}$ and use the exponential decay strategy to adjust the learning rate with a factor of 0.98. The loss coefficient $\lambda_{1}$ is set to 0.5. We split all datasets into the training, validation, and test sets as 7:1:2, and all images are resized to $1024\times 1024$ during the fine-tuning and evaluation stage [45]. The batch size and the training epoch are set to 16 and 300. Our Co-Seg++ introduces an additional 8.3% learnable parameters compared to the baseline (in Table 7).
@@ 4.2 Comparison on Medical Semantic Segmentation
Table 2 demonstrates the effectiveness of Co-Seg++ on dental CBCT semantic segmentation. We observe that general medical segmentation methods (e.g, nnUNet [7]) achieve competitive baseline performance, while recent Mamba and prompt approaches further promote dental segmentation. Remarkably, Co-Seg++ achieves state-of-the-art performance across all anatomical structures with statistical significance (e.g., P-value $<$ 0.005). For lower jawbone segmentation, Co-Seg++ reveals 96.54% Dice and 93.85% mIoU, representing a 1.76% Dice improvement over the best baseline ToothSC-SAM [36]. The improvement is even more pronounced for upper jawbone segmentation, where Co-Seg++ achieves 92.78% Dice with a 2.31% improvement. For pharynx segmentation, Co-Seg++ demonstrates consistent superiority with 97.73% Dice and the lowest HD of 13.27, indicating exceptional precision in boundary delineation. These comparisons validate the superior performance of our Co-Seg++ on diverse medical semantic segmentation tasks through mutual optimization.
@@ 4.3 Comparison on Medical Instance Segmentation
We further evaluate Co-Seg++ on medical instance segmentation tasks to demonstrate the effectiveness of our collaborative learning framework in handling complex instance-level segmentation challenges. Table 3 presents comprehensive results on histopathology instance segmentation across multiple cell types and datasets. We observe that specialized histopathology methods generally outperform generic segmentation approaches. CellPose [2] and PathoSAM [15] demonstrate superior performance compared to traditional methods such as nnUNet [7]. Our Co-Seg++ consistently surpasses all baselines by substantial margins across all datasets and cell categories with a P-value $<0.001$. On the PUMA dataset, Co-Seg++ achieves remarkable performance for tumor instance segmentation with 80.25% F1 score and 70.59% AJI, representing significant improvements of 3.71% F1 and 4.97% AJI over the best baseline PathoSAM [15].
@@ 4.4 Comparison on Medical Panoptic Segmentation
To comprehensively evaluate the versatility of Co-Seg++, we conduct extensive experiments on medical panoptic segmentation, which simultaneously requires both semantic understanding and instance-level discrimination capabilities. Table 5 demonstrates the superior performance of Co-Seg++ on histopathology panoptic segmentation across diverse tissue, gland and nuclei types. We observe that specialized methods like CellViT [2] and PathoSAM [15] achieve competitive performance compared to general segmentation approaches, with PathoSAM [15] showing strong results across most categories. In contrast, Co-Seg++ consistently outperforms all baseline methods with substantial improvements and statistical significance (e.g., with a P-value $<0.001$) across all anatomical structures and datasets. On the PUMA dataset, Co-Seg++ achieves remarkable performance for tumor panoptic quality with 64.52% (2.35% improvement over PathoSAM [15]) and demonstrates particularly significant improvement for the challenging stroma segmentation task with 41.67% (3.22% improvement). Moreover, we compare the quantitative results of our Co-Seg++ with the best baseline methods in Fig. 6 and our UN-SAM can segment nuclei more accurately with better tissue and gland regions.
@@ 4.5 Ablation Study
To investigate the effectiveness of the STP-Encoder $P$, MTC-Decoder $D$, and co-segmentation paradigm $C$, we conduct comprehensive ablation studies on histopathology panoptic segmentation across three datasets (PUMA, CRAG, and GlaS), as illustrated in Table 7. By removing the tailored modules from Co-Seg, we construct an original encoder-shared framework with Hiera-L as the ablation baseline. By separately introducing the STP-Encoder ($2^{nd}$ row), MTC-Decoder ($3^{rd}$ row), and co-segmentation paradigm ($4^{th}$ row), the performance demonstrates consistent improvements across all datasets. Notably, our co-segmentation paradigm achieves the most significant individual contribution, with substantial improvements averaging PQ of 1.61%. We additionally investigate the effects of combined components. By comparing $6^{th}$ and $7^{th}$ rows with the $5^{th}$ row, configurations with co-segmentation demonstrate significant gains, achieving an average PQ of 53.48% and 53.20%, respectively. On this basis, our complete Co-Seg++ framework ($8^{th}$ row) simultaneously adopts all three components to achieve the best performance across all datasets and categories. The full model attains optimal performance with an average PQ of 54.25%, representing an average improvement of 3.10% over the baseline. These comprehensive ablation experiments validate that the tailored STP-Encoder, MTC-Decoder, and co-segmentation paradigm collectively contribute to the superior performance of Co-Seg++ by effectively capturing long-range dependencies, enabling cross-task collaboration, and facilitating mutual enhancement between semantic and instance segmentation.
@@ 5 Conclusion
In this work, we identify the interdependencies between semantic and instance segmentation in medical imaging, and propose a novel co-segmentation paradigm to establish the Co-Seg++ framework, allowing semantic and instance segmentation tasks to mutually enhance each other. The framework integrates an STP-Encoder to capture long-range spatial and temporal relationships as prior constraints, and an MTC-Decoder that leverages cross-guidance to strengthen contextual consistency across tasks. Extensive experiments on diverse histopathology and CBCT datasets demonstrate that Co-Seg++ surpasses state-of-the-art methods across versatile segmentation, validating the effectiveness of our collaborative learning approach for comprehensive medical image understanding.

### v2

@@ 4.1.1 Datasets
[…]
@@ 4.1.2 Implementation Details
We perform all experiments on a workstation equipped with an Intel Xeon Silver 4216 CPU, 256GB RAM, and an NVIDIA A5000 GPU (24GB memory). The software environment consists of PyTorch 2.4 and CUDA 12.8. For fair comparisons, all semantic and instance segmentation methods are implemented with the same training settings and configurations. We utilize the pre-trained SAM ViT-H [32] structure as the image encoder of medical SAM baselines. We adopt the Hirea-L structure [42] as the image encoder of our Co-Seg++. We apply the optimizer using Adam with an initial learning rate of $1\times 10^{-4}$ and use the exponential decay strategy to adjust the learning rate with a factor of 0.98. The loss coefficient $\lambda_{1}$ is set to 0.5. We split all datasets into the training, validation, and test sets as 7:1:2, and all images are resized to $1024\times 1024$ during the fine-tuning and evaluation stage [44]. The batch size and the training epoch are set to 16 and 300. Our Co-Seg++ introduces 8.3% additional learnable parameters compared to the baseline in Table 7.
@@ 4.2 Comparison on Medical Semantic Segmentation
Table 2 demonstrates the effectiveness of Co-Seg++ on dental CBCT semantic segmentation. We observe that general medical segmentation methods (e.g, nnUNet [7]) achieve competitive baseline performance, while recent Mamba and prompt approaches further promote dental segmentation. Remarkably, Co-Seg++ achieves state-of-the-art performance across all anatomical structures with statistical significance (e.g., P-value << 0.005). For lower jawbone segmentation, Co-Seg++ reveals 96.54% Dice and 93.85% mIoU, representing a 1.76% Dice improvement over the best baseline ToothSC-SAM [41]. The improvement is even more pronounced for upper jawbone segmentation, where Co-Seg++ achieves 92.78% Dice with a 2.31% improvement. For pharynx segmentation, Co-Seg++ demonstrates consistent superiority with 97.73% Dice and the lowest HD of 13.27, indicating exceptional precision in boundary delineation. These comparisons validate the superior performance of our Co-Seg++ on diverse medical semantic segmentation tasks through mutual optimization.
@@ 4.3 Comparison on Medical Instance Segmentation
We further evaluate Co-Seg++ on medical instance segmentation tasks to demonstrate the effectiveness of our collaborative learning framework in handling complex instance-level segmentation challenges. Table 3 presents comprehensive results on histopathology instance segmentation across multiple cell types and datasets. We observe that specialized histopathology methods generally outperform generic segmentation approaches. CellPose [2] and PathoSAM [16] demonstrate superior performance compared to traditional methods such as nnUNet [7]. Our Co-Seg++ consistently surpasses all baselines by substantial margins across all datasets and cell categories with a P-value <0.001<0.001. On the PUMA dataset, Co-Seg++ achieves remarkable performance for tumor instance segmentation with 80.25% F1 score and 70.59% AJI, representing significant improvements of 3.71% F1 and 4.97% AJI over the best baseline PathoSAM [16].
@@ 4.4 Comparison on Medical Panoptic Segmentation
To comprehensively evaluate the versatility of Co-Seg++, we conduct extensive experiments on medical panoptic segmentation, which simultaneously requires both semantic understanding and instance-level discrimination capabilities. Table 5 demonstrates the superior performance of Co-Seg++ on histopathology panoptic segmentation across diverse tissue, gland and nuclei types. We observe that specialized methods like CellViT [2] and PathoSAM [16] achieve competitive performance compared to general segmentation approaches, with PathoSAM [16] showing strong results across most categories. In contrast, Co-Seg++ consistently outperforms all baseline methods with substantial improvements and statistical significance (e.g., with a P-value <0.001<0.001) across all anatomical structures and datasets. On the PUMA dataset, Co-Seg++ achieves remarkable performance for tumor panoptic quality with 64.52% (2.35% improvement over PathoSAM [16]) and demonstrates particularly significant improvement for the challenging stroma segmentation task with 41.67% (3.22% improvement). Moreover, we compare the quantitative results of our Co-Seg++ with the best baseline methods in Fig. 6 and our Co-Seg++ can segment nuclei more accurately with better tissue and gland regions.
@@ 4.5 Ablation Study
To investigate the effectiveness of the SSP-Encoder $P$, MTC-Decoder $D$, and co-segmentation paradigm $C$, we conduct comprehensive ablation studies on histopathology panoptic segmentation across three datasets (PUMA, CRAG, and GlaS), as illustrated in Table 7. By removing the tailored modules from Co-Seg, we construct an original encoder-shared framework with Hiera-L [36] as the ablation baseline. By separately introducing the SSP-Encoder ($2^{nd}$ row), MTC-Decoder ($3^{rd}$ row), and co-segmentation paradigm ($4^{th}$ row), the performance demonstrates consistent improvements across all datasets. Notably, our co-segmentation paradigm achieves the most significant individual contribution, with substantial improvements averaging PQ of 1.61%. We additionally investigate the effects of combined components. By comparing $6^{th}$ and $7^{th}$ rows with the $5^{th}$ row, configurations with co-segmentation demonstrate significant gains, achieving an average PQ of 53.48% and 53.20%, respectively. On this basis, our complete Co-Seg++ framework ($8^{th}$ row) simultaneously adopts all three components to achieve the best performance across all datasets and categories. The full model attains optimal performance with an average PQ of 54.25%, representing an average improvement of 3.10% over the baseline. These comprehensive ablation experiments validate that the tailored SSP-Encoder, MTC-Decoder, and co-segmentation paradigm collectively contribute to the superior performance of Co-Seg++ by effectively capturing long-range dependencies, enabling cross-task collaboration, and facilitating mutual enhancement between semantic and instance segmentation.
Furthermore, we investigate the effect of SiLU activation and gating operations (GO) on our Co-Seg++, as illustrated in Table 8. By introducing the SiLU activation function, we observe performance degradation across all categories, with semantic segmentation dropping from 64.52% to 60.17% for tumor regions and from 41.67% to 36.73% for stroma regions. Similarly, instance segmentation performance decreases from 67.37% to 62.15% for tumor nuclei and from 54.60% to 51.71% for lymphocytes. When both SiLU and gating operations are added, the performance shows slight improvement compared to SiLU-only but still remains inferior to our streamlined design. Therefore, the additional complexity introduced by SiLU activation and gating operations is not beneficial for our framework. These results demonstrate that the design of our SSP-Encoder is both effective and efficient, achieving superior performance in versatile medical segmentation.
Table 8: Ablation Study of SSP-Encoder on PUMA panoptic segmentation.
Methods Semantic Instance
Tumor Stroma Tumor Lymph Other Cells
SSP-Encoder + SiLU 60.17 36.73 62.15 51.71 25.46
SSP-Encoder + SiLU + GO 62.03 38.09 64.92 52.43 27.75
SSP-Encoder (Ours) 64.52 41.67 67.37 54.60 29.48
Table 9: Comparison of computation cost on histopathology semantic and instance segmentation.
[…]
Inference
[…]
Params
[…]
Learnable
[…]
CellViT [2] 4896.15 823.68 891.29 891.29 80.61 43.57
PathoSAM [16] 2739.59 448.33 645.16 645.16 81.46 45.53
Co-Seg++ 869.67 186.53 234.33 22.18 84.09 49.56
Figure 8: Hyper-parameter analysis of $\lambda_{1}$ and number of forwards on the PUMA panoptic segmentation. The metrics are averaged PQ across two semantic categories (SC) and three instance categories (IC).
[…]
Table 10: Evaluation on domain shift across different histopathology datasets.
[…]
Methods Tumor Stroma Nucleus Epithelial LR Connective
[…]
nnUnet [7] 26.43 19.15 582.37 23.58 16.42 545.28 41.26 39.13 42.37 34.16 29.54 19.82 18.43 15.67
Swin-Umamba [12] 28.91 21.36 568.24 25.74 18.29 531.65 44.58 42.27 44.82 36.29 31.78 21.54 20.16 17.23
EMCAD [8] 30.17 22.64 561.45 27.35 19.58 524.83 48.35 45.81 46.51 37.94 33.29 22.87 21.74 18.55
H-SAM [3] 32.20 23.63 550.16 29.10 20.88 510.35 63.45 59.39 52.79 41.87 38.64 27.45 27.91 24.25
Zig-RiR [37] 31.48 23.17 556.72 28.26 20.14 518.94 58.87 56.96 49.65 39.83 37.92 26.47 24.38 21.16
HoverNet [9] 34.27 25.18 547.83 30.46 22.15 507.62 56.94 54.23 46.82 37.68 33.51 22.94 24.73 21.18
HookNet [38] 32.65 24.01 555.29 29.38 21.27 516.47 51.82 48.74 43.15 34.82 35.68 24.92 22.57 19.38
CellPose [11] 36.28 27.15 542.68 32.17 23.86 502.91 45.81 46.49 44.29 36.17 35.82 24.58 25.16 21.47
CPPNet [10] 35.19 26.24 546.35 30.59 22.25 501.40 53.75 43.41 50.27 40.18 38.74 27.16 25.31 21.83
CellViT [2] 36.85 27.59 541.27 32.74 24.38 501.15 58.72 55.89 48.13 39.52 34.62 23.71 23.25 20.91
PromptNucSeg [1] 40.56 31.73 552.10 35.92 27.14 498.63 65.92 61.98 54.38 43.17 43.59 30.28 29.47 25.36
PathoSAM [16] 43.72 34.86 525.43 37.15 28.23 489.36 69.27 64.15 55.81 43.92 44.52 31.09 32.18 27.63
Co-Seg++ 51.46 43.39 486.16 40.11 30.61 463.00 74.14 66.23 57.34 44.61 47.01 32.74 34.95 29.38
Table 11: Impact of training data proportion on PUMA panoptic segmentation.
Training Data Methods Semantic Instance
Tumor Stroma Tumor Lymph Other Cells
10% CellViT [2] 42.87 26.14 45.38 35.74 18.52
PromptNucSeg [1] 43.68 26.92 45.96 36.47 19.26
PathoSAM [16] 41.56 25.38 44.17 34.29 17.84
Co-Seg++ 46.05 29.78 48.61 43.48 20.73
30% CellViT [2] 50.16 31.05 53.29 42.81 22.15
PromptNucSeg [1] 51.05 31.58 53.74 43.29 23.08
PathoSAM [16] 49.28 30.16 52.43 41.95 21.37
Co-Seg++ 53.42 34.55 56.73 50.38 14.10
50% CellViT [2] 56.42 34.89 59.18 48.06 24.96
PromptNucSeg [1] 57.12 35.24 59.38 48.67 25.41
PathoSAM [16] 56.89 35.21 60.15 48.52 24.85
Co-Seg++ 59.18 38.26 61.48 53.91 21.73
100% CellViT [2] 61.95 38.28 64.72 52.81 27.69
PromptNucSeg [1] 61.83 38.19 63.58 52.74 27.52
PathoSAM [16] 62.17 38.45 65.01 52.86 27.88
Co-Seg++ 64.52 41.67 67.37 54.60 29.48
4.6 Computation Efficiency Analysis
To evaluate the computation efficiency of Co-Seg++, we conducted comparisons with the independent segmentation model CellViT [2] and the encoder-shared segmentation model PathoSAM [16] in terms of resource costs, including the FLOPs, inference latency, learnable and total parameters, as shown in Table 9. It is important to note that Co-Seg++ and PathoSAM can perform both semantic and instance segmentation through a single unified model, while CellViT [2] is task-specific and requires training two separate models to accomplish both segmentation tasks. Therefore, for CellViT [2], we report the cumulative costs of two models. We observe that our Co-Seg++ demonstrates significant computation efficiency across all metrics. In particular, Co-Seg++ requires 869.67G FLOPs, representing a 5.6× reduction compared to the independent CellViT [2] and 3.2× reduction compared to the encoder-shared PathoSAM [16]. In terms of inference latency, Co-Seg++ achieves 186.53ms per image, which is 4.4× lower than CellViT [2] and 2.4× lower than PathoSAM [16]. Moreover, Co-Seg++ requires only 22.18M learnable parameters, representing a remarkable 40.2× reduction compared to CellViT [2] and 29.1× reduction compared to PathoSAM [16]. Moreover, Co-Seg++ achieves superior performance, outperforming PathoSAM [16] by 2.63% (Dice) and 4.03% (AJI) for semantic and instance segmentation, respectively, across three histopathology datasets. These comprehensive evaluations prove the superior computation efficiency of our Co-Seg++ framework in versatile medical segmentation.
4.7 Domain Shift Assessment
[…]
4.8 Limited Annotation Analysis
To investigate the stability of Co-Seg++ under limited supervision, we conduct sensitivity analysis by training our framework with varying proportions of training data (i.e., 10%, 30%, 50%, and 100%) on the PUMA dataset for panoptic segmentation. As presented in Table 11, Co-Seg++ consistently outperforms state-of-the-art methods across all annotation scales for both semantic and instance segmentation tasks. With only 10% of training data, Co-Seg++ achieves 46.05% and 48.61% panoptic quality for tumor semantic and instance segmentation, representing improvements of 2.37% and 2.65% over PromptNucSeg [1], while for the challenging lymphocyte detection task, Co-Seg++ achieves 43.48% compared to PromptNucSeg's 36.47%, demonstrating a substantial 7.01% improvement. When training data increases to 30% and 50%, Co-Seg++ maintains superiority with consistent improvements of 2-7% across all categories, achieving 53.42% and 59.18% for tumor semantic segmentation at 30% and 50% data respectively, outperforming PromptNucSeg [1] by 2.37% and 2.06%. Even with full annotation availability, Co-Seg++ maintains performance advantages, achieving 64.52% and 41.67% for tumor and stroma semantic segmentation, representing improvements of 2.69% and 3.22% over PathoSAM [16]. Compared to CellViT [2], which employs multi-task learning without collaborative mechanisms, Co-Seg++ demonstrates more stable improvements across all annotation scales, with margins of 3.18% and 7.74% at 10% data and 2.57% and 1.79% at 100% data for tumor semantic and lymphocyte instance segmentation, indicating that the collaborative learning paradigm provides complementary benefits independent of annotation scale. These results demonstrate the robustness of our Co-Seg++ framework in realistic scenarios where comprehensive annotations are challenging to acquire.
Figure 9: Training dynamics comparison of w/ MTC-Decoder and w/o MTC-Decoder on the PUMA dataset with two tasks, including (1) semantic and (2) instance segmentation.
4.9 Interpretability Analysis of Mutual Guidance
To analyze how cross-task guidance in the MTC-Decoder affects prediction dynamics and validate that mutual guidance consistently benefits both tasks, we conduct comprehensive training dynamics analysis and qualitative visualization. We compare two configurations on the PUMA dataset: with (w/) and without (w/o) MTC-Decoder, tracking task-specific performance after each training epoch to examine learning behavior over time. As illustrated in Fig. 9, the configuration w/ MTC-Decoder consistently outperforms w/o MTC-Decoder across all training epochs for both semantic and instance segmentation tasks, with performance improvements of 2-3% maintained from early training stages through convergence. Notably, both tasks benefit from the cross-task guidance mechanism without exhibiting negative transfer, where semantic segmentation achieves 64.52% panoptic quality compared to 61.85% w/o MTC-Decoder, and instance segmentation reaches 67.37% compared to 64.58% w/o MTC-Decoder.
Moreover, to further interpret how mutual guidance improves segmentation results, we perform qualitative visual analysis on the GlaS dataset for gland semantic segmentation and epithelial instance segmentation, as shown in Fig. 10. We visualize the binary mask prompts and cross-attention maps to demonstrate the information flow between tasks. Specifically, the semantic segmentation branch generates region-level prompts that capture global gland structure and contextual information, which effectively guides the instance branch to better distinguish individual cell boundaries within tissue regions by providing spatial constraints. Conversely, the instance segmentation branch produces fine-grained prompts highlighting precise object boundaries and morphological details, which in turn refine the semantic region boundaries and reduce over-segmentation errors. The attention maps reveal that when semantic features act as keys and values while instance features serve as queries, the cross-attention mechanism successfully integrates global contextual information into instance-level features, enabling the model to resolve ambiguous boundaries in densely packed or overlapping nuclei. Similarly, when instance features guide semantic segmentation, the fine-grained boundary information helps establish more precise tissue region delineation. These analyses validate that the MTC-Decoder enables both tasks to mutually enhance each other, where complementary contextual dependencies are effectively captured to improve segmentation consistency and accuracy.
Figure 10: Visualization of cross-task guidance mechanism with binary mask prompts and attention maps on the GlaS dataset. Our Co-Seg++ with MTC-Decoder exhibits superior coherent semantic masks and accurate instance boundaries.
@@ 5 Conclusion
In this work, we identify the interdependencies between semantic and instance segmentation in medical imaging, and propose a novel co-segmentation paradigm to establish the Co-Seg++ framework, allowing semantic and instance segmentation tasks to mutually enhance each other. The framework integrates a SSP-Encoder to capture long-range spatial and temporal relationships as prior constraints, and a MTC-Decoder that leverages cross-guidance to strengthen contextual consistency across tasks. Extensive experiments on diverse histopathology and CBCT datasets demonstrate that Co-Seg++ surpasses state-of-the-art methods across versatile segmentation, validating the effectiveness of our collaborative learning approach for comprehensive medical image understanding. Future work will focus on extending the co-segmentation paradigm to 3D volumetric medical image analysis, such as CT and MRI segmentation, and exploring its applicability to multi-modal medical imaging tasks.